# SpotV2Net: Multivariate Intraday Spot Volatility Forecasting via Vol-of-Vol-Informed Graph Attention Networks

### v2

@@ Abstract
50127 Firenze FI (Italy). E-mail: giacomo.toscano@unifi.it
@@ 1 Introduction
In a multidimensional setting, we assume that each asset represents a node of a graph. The features (covariates) associated with each node comprise the instantaneous (or spot) volatility of the asset and the instantaneous co-volatilities with every other asset in the graph, reconstructed using the non-parametric Fourier methodology by Malliavin and Mancino, (2002, 2009). Further, to capture spillover effects, the edge features encapsulate information about the dynamic dependence between node pairs. Specifically, for any given edge, its features consist of Fourier estimates of the volatility of volatility of each of the assets connected by the edge and the co-volatility of their individual volatilities. In this regard, we note that different studies have shown that encompassing realized or implied volatility of volatility measures improves a model’s performance in risk management tasks (see, e.g., Huang et al., (2019); Chen et al., (2021); Li, (2022) and forecasting applications (see, e.g., Catania and Proietti, (2020); Campisi et al., (2023); Ding, (2023)). Node and edge features represent the inputs of the forecasting model. To account for (cross) serial dependence, each input includes not only current values but also several lagged values. This is achieved by leveraging the capability of deep learning architectures to handle a large number of inputs (Jaegle et al.,, 2021; Menghani,, 2023). In the specific case of GNNs, inputs can be placed both in the nodes and the edges (Wu et al., 2020a, ; Zhou et al.,, 2020). We refer to the resulting architecture that we obtain as Spot Volatility and Volatility of volatility Network, abbreviated to SpotV2Net.
@@ 2 Related literature
Within a multivariate framework, forecasting an intraday spot measure of asset volatility is a critical area of exploration that has received little attention in the literature, to the best of our knowledge. Few works attempt to model intraday volatility through ARCH (Taylor and Xu,, 1997) and GARCH (Engle and Sokalska,, 2012) variants. However, their modeling focus is on integrated measures of volatilities, namely on realized volatilities (Andersen et al.,, 2001), rather than on spot estimates. Moreover, while these models can capture linear intraday patterns through multiplicative components (Engle and Sokalska,, 2012), they may not fully capture more complex nonlinear intraday patterns that characterize the microstructure of financial markets.
On this behalf, the usage of ML techniques in finance has recently marked an impact on the field. The capability of ML models, especially deep neural networks, to handle high-dimensional data and approximate complex patterns has impacted key areas of finance, such as portfolio optimization (Heaton et al.,, 2017; Zhang et al.,, 2020; Ma et al.,, 2021; Lin and Taamouti,, 2023), asset pricing (Gu et al.,, 2020; Wu et al.,, 2021; Chen et al.,, 2023) and volatility forecasting (Xiong et al.,, 2015; Liu,, 2019; Bucci,, 2020; Christensen et al.,, 2022; Zhu et al.,, 2023). For the specific problem of volatility forecasting, it is worth underlying that the findings by Christensen et al., (2022) suggest that several ML algorithms, despite being implemented with minimal hyperparameter tuning, yield improved forecasting performance compared to models in the HAR lineage, especially at longer horizons. The authors attribute such forecast accuracy gains to higher persistence in the ML models - which helps to approximate the long memory of realized variance - and the capability of ML models to encompass the contribution from additional predictors, besides past volatility lags.
@@ 4 Motivation
Figure 1: Fourier spot volatility estimates for AXP, CAT and V, obtained every 30 minutes using 1-second prices. Details about the Fourier estimator of the spot volatility are provided in Sec. 6.
However, by observing Fig. 2, one immediately notices that the pairwise co-volatilities of volatility behave very differently. In particular, the spot co-volatility of volatility between AXP and V shows two very large peaks near March 16 and March 17, while the co-volatility of volatility between each of these two stocks and CAT maintains substantially stable and lower values during March 2023. Such bursts in the spot co-volatility of volatility may indicate the presence of more pronounced instantaneous spillover effects between AXP and V, compared to the pairs AXP-CAT and V-CAT. This is not surprising, since AXP and V operate in the same sector, namely, financial services, which the collapse of Silicon Valley Bank and Signature Bank directly impacted. Instead, CAT is a well-diversified manufacturing company whose spot volatility may be less correlated with the volatility of companies operating in the banking and insurance industry. SpotV2Net can recognize such effects using the attention mechanism inherent in the GAT architecture. When predicting the future value of a specific node, the attention mechanism allows assigning different weights to neighbor nodes, based not only on node features but also on features embedded in the connecting edges. Hence, the attention mechanism enables SpotV2Net to naturally capture the time-varying spillover effects between volatility pairs by including edge features. The next Section describes the attention mechanism with detail.
Figure 2: Fourier spot co-volatility of volatility estimates for the pairs AXP-CAT, AXP-V and CAT-V, obtained every 30 minutes using 1-second prices. Details about the Fourier estimator of the spot co-volatility of volatility are provided in Sec. 6.
@@ 5 The SpotV2Net Model
Let us now consider the mechanism behind a GAT. Let $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ denote a graph structure, where $\mathcal{V}=\left\{\nu_{1},\ldots,\nu_{N}\right\}$ is the set of the $N$ nodes of the structure and $\mathcal{E}$ is the set of edges, where $\varepsilon_{ij}=\left(\nu_{i},\nu_{j}\right)\in\mathcal{E}$ denotes the edge connecting node $\nu_{i}$ and node $\nu_{j}$. Further, let ${A}$ denote the $N\times N$ binary adjacency matrix of the structure, where ${A}[i,j]$ is equal to $1$ (respectively, 00) if $\nu_{i}$ and $\nu_{j}$ are connected (respectively, are not connected). The graph $\mathcal{G}$ is assumed to be fully connected, that is, all the entries of $A$ are assumed to be equal to $1$.
@@
The input to the first GAT layer is the set of vectors representing node features, namely $x=\{x_{1},x_{2},\ldots,x_{N}\}$, $x_{i}\in\mathbb{R}^{M}$, where $M$ indicates the number of features in each node. The GAT layer outputs a different representation of the initial set of node features, with different cardinality $M^{\prime}$, denoted by $x^{\prime}=\{x^{\prime}_{1},x^{\prime}_{2},\ldots,x^{\prime}_{N}\}$, $x^{\prime}_{i}\in\mathbb{R}^{M^{\prime}}$, see Fig. 3. We treat $M^{\prime}$ as a hyperparameter and optimize its value through the training phase.
Figure 3: Transformation of node features through a Graph Attentional Layer. Initially, nodes $x_{1},x_{2},\ldots,x_{6}$ possess individual features. Then, the attentional layer selectively aggregates information from neighboring nodes, factoring in their relative importance through learned attention weights. The result is a set of updated node features $x^{\prime}_{1},x^{\prime}_{2},\ldots,x^{\prime}_{6}$, each reflecting the combined influence of its neighbors within the graph.
@@
$x_{i}=\left[\left\{\widehat{V}_{i,b-l}\right\}_{l=0,\ldots,L}\,,\,\left\{\left\{\widehat{C}_{ij,b-l}\right\}_{j=1,\ldots,N;j\neq i}\right\}_{l=0,1,\ldots,L}\right],$
@@
$x_{ij}^{e}=\left[\left\{\widehat{\widetilde{V}}_{i,b-l}\right\}_{l=0,\ldots,L}\,,\,\left\{\widehat{\widetilde{V}}_{j,b-l}\right\}_{l=0,\ldots,L}\,,\,\left\{\widehat{\widetilde{C}}_{ij,b-l}\right\}_{l=0,1,\ldots,L}\right],$
@@
Figure 4: Three-dimensional example of the fully connected graph topology of SpotV2Net. The three nodes are denoted by $i$,$j$ and $k$.
@@ 6 Fourier estimators
The Fourier non-parametric estimation method, originally introduced by Malliavin and Mancino, (2002, 2009), allows reconstructing the latent spot co-volatility paths from high-frequency price observations. The method is particularly apt for estimating spot co-volatilities in the presence of microstructure noise, as it removes the influence of the latter by simply cutting off the highest coefficients of the price increment from the convolution that yields the co-volatility coefficients (see Eq. (1)), without the need of a bias correction.
@@
The Fourier method also allows us to estimate the co-volatility of volatility path (Barucci and Mancino,, 2010). It is worth noting that the Fourier spot co-volatility of volatility estimator does not require the pre-estimation of the spot volatility path but only the computation of integrated quantities, namely the coefficients of the co-volatility. This feature enhances the finite-sample efficiency of the Fourier methodology, compared to alternative approaches that require the pre-estimation of the spot volatility path (see the discussion in Toscano et al., (2022)).
The consistency of the Fourier spot volatility of volatility estimator with noisy prices is proven in Mancino et al., (2022). Its finite-sample performance at high-frequencies is studied in Toscano, (2022).
@@ 6.1 Fourier estimator of the volatility and co-volatility
Without loss of generality, consider the case with just two assets, that is $N=2$. For any $i=1,2$, assume that the log-price process $p_{i}$ is observable on the grid $\mathcal{D}_{n}:=\{0=t_{0,n}<t_{1,n}\ldots<t_{n-1,n}<h=t_{n,n}\}$ such that $max_{s\in\{1,\ldots,n\}}|t_{s,n}-t_{s-1,n}|\to 0$ as $n\to\infty$ [^7] For simplicity of the exposition, we assume that two price series are synchronous, that is, are observed on the same grid $\mathcal{D}_{n}$, containing $n+1$ points..
@@
$c_{k}\left(C_{n,N_{c}}\right)=\frac{h}{2N_{c}+1}\sum_{|l|\leq N_{c}}c_{l}\left(dp_{1,n}\right)c_{k-l}\left(dp_{2,n}\right),$ (1)
@@
$c_{k}\left(dp_{i,n}\right):=\frac{1}{h}\sum_{l=0}^{{n}-1}e^{-\mathrm{i}k\frac{2\pi}{h}t_{l,n}}\left(p_{i}(t_{l+1,n})-p_{i}(t_{l,n})\right),\quad i=1,2.$
@@
$\widehat{C}_{12,b}=\sum_{|k|<M_{c}}\left(1-\frac{|k|}{M_{c}}\right)c_{k}\left(C_{n,N_{c}}\right)e^{\mathrm{i}k\frac{2\pi}{h}\tau_{b}},$
@@
$\widehat{V}_{i,b}=\sum_{|k|<M_{v_{i}}}\left(1-\frac{|k|}{M_{v_{i}}}\right)c_{k}\left(V_{n,N_{v_{i}}}\right)e^{\mathrm{i}k\frac{2\pi}{h}\tau_{b}},\quad i=1,2,$
@@ 6.2 Fourier estimator of the volatility of volatility and the co-volatility of volatility
$c_{k}\left(\widetilde{C}_{n,N_{c},S_{c}}\right)=\frac{h}{2S_{c}+1}\sum_{|l|\leq S_{c}}l(l-k)c_{l}\left(C_{n,N_{c}}\right)c_{k-l}\left(C_{n,N_{c}}\right),$
@@
$c_{k}\left(\widetilde{V}_{n,N_{v_{i}},S_{v_{i}}}\right)=\frac{h}{2S_{v_{i}}+1}\sum_{|l|\leq S_{v_{i}}}l(l-k)c_{l}\left(V_{n,N_{v_{i}}}\right)c_{k-l}\left(V_{n,N_{v_{i}}}\right),\quad i=1,2,$
@@
$\widehat{\widetilde{C}}_{12,b}=\sum_{|k|<L_{c}}\left(1-\frac{|k|}{L_{c}}\right)c_{k}\left(\widetilde{C}_{n,N_{c},S_{c}}\right)e^{\mathrm{i}k\frac{2\pi}{h}\tau_{b}},$
@@
$\widehat{\widetilde{V}}_{i,b}=\sum_{|k|<L_{v_{i}}}\left(1-\frac{|k|}{L_{v_{i}}}\right)c_{k}\left(\widetilde{V}_{n,N_{v_{i}},S_{v_{i}}}\right)e^{\mathrm{i}k\frac{2\pi}{h}\tau_{b}},\quad i=1,2,$
@@ 7 Empirical study: an application to the universe of DJIA stocks
For the implementation of Fourier estimator, we proceed as follows. Starting from tick data, we filter the transactions to include only those executed on the New York Stock Exchange (NYSE) and exclude transactions that occurred outside the operating market hours from 9:30 a.m. to 4:00 p.m. Then, we resample the tick data to the one-second frequency, selecting the price associated with the transaction closest to each second’s onset. The resulting sample size employed in the implementation of the estimators is $n=23400$ per day. The cutting frequencies $N_{v_{i}}$, $M_{v_{i}}$, $S_{v_{i}}$, $L_{v_{i}}$ (for the $i$-th asset) and $N_{c}$, $M_{c}$, $S_{c}$ and $L_{c}$ (for each pair of assets) are selected using the guidance from the paper by Sanfelici and Toscano, (2024), which introduces a MATLAB library for the implementation of Fourier estimators [^8] The Fourier method is not consistent in the presence of jumps. To remove the effect of jumps from the estimation, we identify 1-second returns whose absolute value is larger than a given threshold $\vartheta_{n}$ and replace them with zeros. Letting $\vartheta_{n}=\beta\left(T/n\right)^{\alpha}$, the constants that determines the threshold are selected using numerical simulations. As a result, we find that suitable choices are $\beta=0.5$ and $\alpha=0.5$.. Univariate and multivariate spot volatility and volatility of volatility estimates are obtained on the 30-minute grid.
[…]

### v3

@@ Abstract
Firenze, 50127 (Italy). E-mail: giacomo.toscano@unifi.it
@@ 1 Introduction
In a multidimensional setting, we assume that each asset represents a node of a graph. The features (covariates) associated with each node comprise the instantaneous (or spot) volatility of the asset and the instantaneous co-volatilities with every other asset in the graph, reconstructed using the non-parametric Fourier methodology by Malliavin and Mancino, (2002, 2009). Furthermore, to capture spillover effects, the edge features encapsulate information about the dynamic dependence between node pairs. Specifically, for any given edge, its features consist of Fourier estimates of the volatility of volatility of each of the assets connected by the edge and the co-volatility of their individual volatilities. In this regard, we note that different studies have shown that encompassing realized or implied volatility of volatility measures improves a model’s performance in risk management tasks (see, e.g., Huang et al., (2019); Chen et al., (2021); Li, (2022) and forecasting applications (see, e.g., Catania and Proietti, (2020); Campisi et al., (2023); Ding, (2023)). Node and edge features represent the inputs of the forecasting model. To account for (cross) serial dependence, each input includes not only current values but also several lagged values. This is achieved by leveraging the capability of deep learning architectures to handle a large number of inputs (Jaegle et al.,, 2021; Menghani,, 2023). In the specific case of GNNs, inputs can be placed both in the nodes and the edges (Wu et al., 2020a, ; Zhou et al.,, 2020). We refer to the resulting architecture that we obtain as Spot Volatility and Volatility of volatility Network, abbreviated to SpotV2Net.
@@ 2 Related literature
Within a multivariate framework, forecasting an intraday spot measure of asset volatility is a critical area of exploration that has received little attention in the literature, to the best of our knowledge. Few works attempt to model intraday volatility through ARCH (Taylor and Xu,, 1997) and GARCH (Engle and Sokalska,, 2012) variants. However, their modeling focus is on integrated measures of volatilities, namely on realized volatilities (Andersen et al.,, 2001), rather than on spot estimates. Moreover, as pointed out by Christensen et al., (2023), traditional linear models may break down when the explanatory variables are strongly correlated, exhibit low signal-to-noise ratios, or if the underlying structure is nonlinear.
On this behalf, ML techniques, which have recently made an impact on the field of finance, represent an alternative approach to effectively handling these issues. The capability of ML models, especially deep neural networks, to handle high-dimensional data and approximate complex patterns has impacted key areas of finance, such as portfolio optimization (Heaton et al.,, 2017; Zhang et al.,, 2020; Ma et al.,, 2021; Lin and Taamouti,, 2023), asset pricing (Gu et al.,, 2020; Wu et al.,, 2021; Chen et al.,, 2023) and volatility forecasting (Xiong et al.,, 2015; Liu,, 2019; Bucci,, 2020; Christensen et al.,, 2023; Zhu et al.,, 2023). For the specific problem of volatility forecasting, it is worth underlying that the findings by Christensen et al., (2023) suggest that several ML algorithms, despite being implemented with minimal hyperparameter tuning, yield improved forecasting performance compared to models in the HAR lineage, especially at longer horizons. The authors attribute such forecast accuracy gains to higher persistence in the ML models - which helps to approximate the long memory of realized variance - and the capability of ML models to encompass the contribution from additional predictors, besides past volatility lags.
@@ 4 Motivation
However, by observing Fig. 2, one immediately notices that the pairwise co-volatilities of volatility behave very differently. In particular, the spot co-volatility of volatility between AXP and V shows two very large peaks near March 16 and March 17, while the co-volatility of volatility between each of these two stocks and CAT maintains substantially stable and lower values during March 2023. Such bursts in the spot co-volatility of volatility may indicate the presence of more pronounced instantaneous spillover effects between AXP and V, compared to the pairs AXP-CAT and V-CAT. This is not surprising, since AXP and V operate in the same sector, namely, financial services, which the collapse of Silicon Valley Bank and Signature Bank directly impacted. Instead, CAT is a well-diversified manufacturing company whose spot volatility may be less correlated with the volatility of companies operating in the banking and insurance industry. SpotV2Net can recognize such effects using the attention mechanism inherent in the GAT architecture. When predicting the future value of a specific node, the attention mechanism allows assigning different weights to neighbor nodes, based not only on node features but also on features embedded in the connecting edges. Hence, the attention mechanism enables SpotV2Net to naturally capture the time-varying spillover effects between volatility pairs by including edge features. The next Section describes the attention mechanism in detail.
Figure 1: Fourier spot volatility estimates $\{V_{i,t}\}_{t\in\mathcal{T}}$ for $i\in\{AXP,CAT,V\}$, on the 30-minute grid $\mathcal{T}=\{\tau_{j}=jT/13:\,j=0,1,...,13\}$, where $T$ is a 6.5-hour trading day. Estimates were obtained using 1-second prices. Details about the Fourier estimator of the spot volatility are provided in Sec. 6.
Figure 2: Fourier spot co-volatility of volatility estimates $\{\widetilde{C}_{ij,t}\}_{t\in\mathcal{T}}$ for $i,j\in\{AXP,CAT,V\},\,i\neq j$, on the 30-minute grid $\mathcal{T}=\{\tau_{j}=jT/13:\,j=0,1,...,13\}$, where $T$ is a 6.5-hour trading day. Estimates were obtained using 1-second prices. Details about the Fourier estimator of the spot volatility are provided in Sec. 6.
@@ 5 The SpotV2Net Model
Let us now consider the mechanism behind a GAT. Let $\mathcal{G}=\{\mathcal{V},\mathcal{E}\}$ denote a graph structure, where $\mathcal{V}=\left\{\nu_{1},\ldots,\nu_{N}\right\}$ is the set of the $N$ nodes of the structure and $\mathcal{E}$ is the set of edges, where $\varepsilon_{ij}=\left(\nu_{i},\nu_{j}\right)\in\mathcal{E}$ denotes the edge connecting node $\nu_{i}$ and node $\nu_{j}$. Besides, let ${A}$ denote the $N\times N$ binary adjacency matrix of the structure, where ${A}[i,j]$ is equal to $1$ (respectively, 00) if $\nu_{i}$ and $\nu_{j}$ are connected (respectively, are not connected). The graph $\mathcal{G}$ is assumed to be fully connected, that is, all the entries of $A$ are assumed to be equal to $1$.
@@
The input to the first GAT layer is the set of vectors representing node features, namely $x=\{x_{1},x_{2},\ldots,x_{N}\}$, $x_{i}\in\mathbb{R}^{M}$, where $M$ indicates the number of features in each node. The GAT layer outputs a different representation of the initial set of node features, with different cardinality $M^{\prime}$, denoted by $x^{\prime}=\{x^{\prime}_{1},x^{\prime}_{2},\ldots,x^{\prime}_{N}\}$, $x^{\prime}_{i}\in\mathbb{R}^{M^{\prime}}$. We treat $M^{\prime}$ as a hyperparameter and optimize its value through the training phase.
@@
$x_{i}=\left[\left\{\widehat{V}_{i,b-l}\right\}_{l=0,\ldots,L}\,,\,\left\{\left\{\widehat{C}_{ij,b-l}\right\}_{j=1,\ldots,N;j\neq i}\right\}_{l=0,1,\ldots,L}\right],$ (1)
@@
$x_{ij}^{e}=\left[\left\{\widehat{\widetilde{V}}_{i,b-l}\right\}_{l=0,\ldots,L}\,,\,\left\{\widehat{\widetilde{V}}_{j,b-l}\right\}_{l=0,\ldots,L}\,,\,\left\{\widehat{\widetilde{C}}_{ij,b-l}\right\}_{l=0,1,\ldots,L}\right],$ (2)
@@
[…]
@@ 6 Fourier estimators
In this Section, we briefly illustrate the Fourier estimation method, which allows us to efficiently reconstruct the entries of node and edge feature vectors defined, respectively, in (1) and (2). The Fourier non-parametric estimation method, originally introduced by Malliavin and Mancino, (2002, 2009), allows reconstructing the latent spot co-volatility paths from high-frequency price observations. The method is particularly apt for estimating spot co-volatilities in the presence of microstructure noise, as it removes the influence of the latter by simply cutting off the highest coefficients of the price increment from the convolution that yields the co-volatility coefficients (see Eq. (3)), without the need of a bias correction.
@@
The Fourier method also allows us to estimate the co-volatility of volatility path (Barucci and Mancino,, 2010). It is worth noting that the Fourier spot co-volatility of volatility estimator does not require the pre-estimation of the spot volatility path but only the computation of integrated quantities, namely the coefficients of the co-volatility. This feature enhances the finite-sample efficiency of the Fourier methodology, compared to alternative approaches that require the pre-estimation of the spot volatility path (see Toscano et al., (2022)). The consistency of the Fourier spot volatility of volatility estimator with noisy prices is proven in Mancino et al., (2022). Its finite-sample performance at high frequencies is studied in Toscano, (2022).
@@ 6.1 Fourier estimator of the volatility and co-volatility
Without loss of generality, consider the case with just two assets, that is $N=2$. For any $i=1,2$, assume that the log-price process $p_{i}$ is observable on the grid $\mathcal{D}_{n}:=\{0=t_{0,n}<t_{1,n}\ldots<t_{n-1,n}<T=t_{n,n}\}$ such that $max_{s\in\{1,\ldots,n\}}|t_{s,n}-t_{s-1,n}|\to 0$ as $n\to\infty$ [^7] For simplicity of the exposition, we assume that two price series are synchronous, that is, are observed on the same grid $\mathcal{D}_{n}$, containing $n+1$ points..
@@
$c_{k}\left(C_{n,N_{c}}\right)=\frac{T}{2N_{c}+1}\sum_{|l|\leq N_{c}}c_{l}\left(dp_{1,n}\right)c_{k-l}\left(dp_{2,n}\right),$ (3)
@@
$c_{k}\left(dp_{i,n}\right):=\frac{1}{T}\sum_{l=0}^{{n}-1}e^{-\mathrm{i}k\frac{2\pi}{T}t_{l,n}}\left(p_{i}(t_{l+1,n})-p_{i}(t_{l,n})\right),\quad i=1,2.$
@@
$\widehat{C}_{12,b}=\sum_{|k|<M_{c}}\left(1-\frac{|k|}{M_{c}}\right)c_{k}\left(C_{n,N_{c}}\right)e^{\mathrm{i}k\frac{2\pi}{T}\tau_{b}},$
@@
$\widehat{V}_{i,b}=\sum_{|k|<M_{v_{i}}}\left(1-\frac{|k|}{M_{v_{i}}}\right)c_{k}\left(V_{n,N_{v_{i}}}\right)e^{\mathrm{i}k\frac{2\pi}{T}\tau_{b}},\quad i=1,2,$
@@ 6.2 Fourier estimator of the volatility of volatility and the co-volatility of volatility
$c_{k}\left(\widetilde{C}_{n,N_{c},S_{c}}\right)=\frac{T}{2S_{c}+1}\sum_{|l|\leq S_{c}}l(l-k)c_{l}\left(C_{n,N_{c}}\right)c_{k-l}\left(C_{n,N_{c}}\right),$
@@
$c_{k}\left(\widetilde{V}_{n,N_{v_{i}},S_{v_{i}}}\right)=\frac{T}{2S_{v_{i}}+1}\sum_{|l|\leq S_{v_{i}}}l(l-k)c_{l}\left(V_{n,N_{v_{i}}}\right)c_{k-l}\left(V_{n,N_{v_{i}}}\right),\quad i=1,2,$
@@
$\widehat{\widetilde{C}}_{12,b}=\sum_{|k|<L_{c}}\left(1-\frac{|k|}{L_{c}}\right)c_{k}\left(\widetilde{C}_{n,N_{c},S_{c}}\right)e^{\mathrm{i}k\frac{2\pi}{T}\tau_{b}},$
@@
$\widehat{\widetilde{V}}_{i,b}=\sum_{|k|<L_{v_{i}}}\left(1-\frac{|k|}{L_{v_{i}}}\right)c_{k}\left(\widetilde{V}_{n,N_{v_{i}},S_{v_{i}}}\right)e^{\mathrm{i}k\frac{2\pi}{T}\tau_{b}},\quad i=1,2,$
@@ 7 Empirical study: an application to the universe of DJIA stocks
For the implementation of Fourier estimator, we proceed as follows. Starting from tick data, we filter the transactions to include only those executed on the New York Stock Exchange (NYSE) and exclude transactions that occurred outside the operating market hours from 9:30 a.m. to 4:00 p.m. Then, we resample the tick data to the one-second frequency, selecting the price associated with the transaction closest to each second’s onset. The resulting sample size employed in the implementation of the estimators is $n=23400$ per day. Note that we measure time in days and set T equal to 1. The cutting frequencies $N_{v_{i}}$, $M_{v_{i}}$, $S_{v_{i}}$, $L_{v_{i}}$ (for the $i$-th asset) and $N_{c}$, $M_{c}$, $S_{c}$ and $L_{c}$ (for each pair of assets) are selected using the guidance from the paper by Sanfelici and Toscano, (2024), which introduces a MATLAB library for the implementation of Fourier estimators [^8] The Fourier method is not consistent in the presence of jumps. To remove the effect of jumps from the estimation, we identify 1-second returns whose absolute value is larger than a given threshold $\vartheta_{n}$ and replace them with zeros. Letting $\vartheta_{n}=\beta\left(T/n\right)^{\alpha}$, the constants that determines the threshold are selected using numerical simulations. As a result, we find that suitable choices are $\beta=0.5$ and $\alpha=0.5$.. Univariate and multivariate spot volatility and volatility of volatility estimates are obtained on the 30-minute grid.
As the output of the univariate estimation process, for each of the $N=30$ DJIA components, we obtain a time series that contains 10318103181031810318999The dataset covers $737$ days. For each day (i.e., $6.5$ hours) we obtain $14$ estimates on the 30-minute grid (note that the last estimate was performed at 3:59 p.m. instead of 4:00 p.m., to reduce the effects of periodicity on Fourier estimates). Hence the total of $737\times 14=10318$ estimates. spot volatility values and a time series with the same length that contains spot volatility of volatility values. Furthermore, the output of the multivariate estimation process includes $\frac{1}{2}N\times(N-1)$ time series with length $10318$ for the spot co-volatilities and other $\frac{1}{2}N\times(N-1)$ time series of the same length for the spot co-volatilities of volatility.